OpenStudy (lgbasallote):

honestly i havent tried solving this yet...but i do need to see how it is solved as i still do NOT know how to start this :/ \[\Large \int_{0}^{1} \frac{\ln x}{\sqrt x} dx\]

OpenStudy (anonymous):

This problem scares me....

OpenStudy (lgbasallote):

i am thinking integration by parts?

OpenStudy (anonymous):

sorta all of it *gulp*

OpenStudy (lgbasallote):

btw...for the record i did not get how to solve my previous question -_- if anyone has an idea do check http://openstudy.com/users/lgbasallote#/updates/4fa872e9e4b059b524f46f3a

OpenStudy (lgbasallote):

^i am troubled by the last part during the plugging in of limits anyway..back tot his one...

OpenStudy (lgbasallote):

@Mimi_x3 ? :(

OpenStudy (mimi_x3):

Integration by parts. u =lnx dv/sqrt(x) du=1/x , u=2sqrt(x)

OpenStudy (lgbasallote):

ugh...i gotta eat -__- owl be back in 30 minutes :/

OpenStudy (mimi_x3):

i made a typo. u=lnx dv=1/sqrt(x) du=1/x , v = 2sqrt(x)

OpenStudy (mimi_x3):

because i feel bad not completing it so.. \[\int\limits\frac{lnx}{\sqrt{x}} dx => \int\limits lnx*\frac{1}{\sqrt{x}} dx\] \(u=lnx\) => dv=1/sqrt(x) => v = 2sqrt(x) \(du= \frac{1}{x}dx\) \[=>lnx*2\sqrt{x} -\int\limits2\sqrt{x}*\frac{1}{x} dx => lnx*2\sqrt{x} -2\int\frac{\sqrt{x}}{x}dx\] you can do it now i think..

OpenStudy (lgbasallote):

my biggest problem in improper integrals is the plugging in NOT the integration -_-

OpenStudy (mimi_x3):

this does not seem like improper integrals..

OpenStudy (mimi_x3):

since the limit is 1 to 0 not infinity to 0

OpenStudy (lgbasallote):

well zero makes it divergent :p

OpenStudy (mimi_x3):

wait yeah sorry i didnt read the question properly; sorry i cant help you. xD

OpenStudy (lgbasallote):

awww :(

OpenStudy (anonymous):

tke lnx =t

OpenStudy (lgbasallote):

uh huh?

OpenStudy (anonymous):

dt=dx/x x=exp(t)

OpenStudy (lgbasallote):

exp (t)?

OpenStudy (anonymous):

exponential of t \[e ^{t}\]

OpenStudy (lgbasallote):

uhmm uhmm i think IPB works better...i just wanna know the steps AFTER the integration

OpenStudy (anonymous):

whats IPB

OpenStudy (lgbasallote):

integration by parts

OpenStudy (lgbasallote):

IBP sorry :))

OpenStudy (anonymous):

the equation becomes \[\int\limits_{a}^{b} te ^{t/2}dt\]

OpenStudy (anonymous):

now it is easy to solve

OpenStudy (anonymous):

use IBP

OpenStudy (lgbasallote):

uhmm..like i said...what i need to know is AFTER the integration...when plugging in those limits..that's where i get confused -_-

OpenStudy (blockcolder):

You use L'hospital's rule.

OpenStudy (lgbasallote):

huh? pls explain further :P

OpenStudy (blockcolder):

Assuming mimi's integration is correct, the answer should be: \[\lim_{t \rightarrow 0^+}(\ln{x}\sqrt{x}-4\sqrt{x})|_t^1\\ =-\sqrt{4}-\lim_{t \rightarrow 0^+}(\ln{t}\sqrt{t}-4\sqrt{t})\] Since \(\lim_{t \rightarrow 0^+}\ln{t}=-\infty\) while \(\lim_{t \rightarrow 0^+}\sqrt{t}=0\), there's an indeterminate form 0*infinity. So you use L'hospital's rule on \(\large \lim_{t \rightarrow 0^+}\frac{\ln{t}}{t^{-1/2}}\).

OpenStudy (lgbasallote):

i see`

OpenStudy (experimentx):

\[ \lim_{t \rightarrow 0^+}(\ln{x}\sqrt{x}-4\sqrt{x})|_t^1\ =-4-\lim_{t \rightarrow 0^+}(\ln{t}\sqrt{t}-4\sqrt{t}) \]

OpenStudy (blockcolder):

Oops. My bad. Stupid typo. =))

OpenStudy (lgbasallote):

i see..and 4 sqrt t is 0 right?

OpenStudy (experimentx):

y

OpenStudy (lgbasallote):

4(0) is zero right o.O